# Simultaneous Task Allocation and Planning for Multi-Robots under Hierarchical Temporal Logic Specifications

### v3

@@ 20            $\Phi^{acc}=\Phi^{acc}\cup\Phi^{acc,k}\,\cup\,\texttt{Children}(\Phi^{acc,k})$ ;
Every sc-LTL formula can be transformed into a Deterministic Finite Automaton (DFA). We opt for NBA, a form of Nondeterministic Finite Automaton (NFA), because sc-LTLs are transformable into either DFAs or NFAs, making both equally expressive for sc-LTL in this sense. The work from [8] and the temporal order heuristic by [21] that our approach in Sec. VI is based on, uses NFAs. Nonetheless, our approach can easily be adapted to work with DFAs as well, as graphically speaking, NFAs are more complex than DFAs.
@@
The syntax and semantics for hierarchical sc-LTL here are designed for robotics and may not be applicable to other domains. Our experience indicates that a top-down reasoning approach is effective in constructing hierarchical sc-LTL specifications from task descriptions. This entails recognizing the inclusion relationships among tasks and organizing them into various levels of abstraction, ranging from general to specific. At each level, tasks that are closely related should be combined into single formulas, while those with weaker connections should be grouped into separate formulas.
@@ IV-C Expressiveness
Temporal logics are interpreted over paths in finite Kripke structures. The Kripke structure is identical to the transition system as presented in Def. III.2 in this work. A Kripke structure, denoted by ${\mathcal{K}}$, is deemed finite when it has a finite number of states. ${\mathcal{K}}$ is considered a model for a temporal logic formula $\phi$ if every initial infinite path within ${\mathcal{K}}$ satisfies $\phi$.
@@
Hierarchical sc-LTL is more expressive than the flat sc-LTL.
@@
We first show that, for any flat sc-LTL, there exists an equivalent hierarchical counterpart. Therefore, the flat sc-LTL is subsumed by the hierarchical sc-LTL. Given a flat sc-LTL $\phi$, an equivalent two-level hierarchical sc-LTL can be constructed in a straightforward manner, where the top level is $\Diamond\phi$ and the bottom level is $\phi$ itself, that is,
@@
Next, we introduce a counter-example to demonstrate the absence of the flat equivalent for certain hierarchical sc-LTL specifications. Considering a Kripke structure with two atomic propositions $\{a,b\}$, the following hierarchical sc-LTL specifications include two mutually exclusive leaf specifications:
@@
$L_{2}:\quad$ $\phi_{2}^{1}=\neg b\,\mathcal{U}\,(a\wedge\neg b)$
$\phi_{2}^{2}=\neg a\,\mathcal{U}\,(b\wedge\neg a).$
While leaf specifications are mutually exclusive, the introduction of the state-specification plan enables the association of each state with one specification, facilitating the separate treatment of leaf specifications. For example, when these two specifications are assigned to separate robots, they can be realistically fulfilled. However, when solely using atomic propositions $a$ and $b$, combining two leaf specifications into a single sc-LTL specification is unfeasible, as it inevitably results in a contradiction, reducing the specification to False. ∎
Remark IV.11
Hierarchical sc-LTL retains the expressive capability of sc-LTL for defining robot tasks, while also enhancing the ability to specify tasks for multiple robots. Moreover, as detailed in Section VIII-A, we conducted a user study that showed hierarchical sc-LTL is more user-friendly than flat sc-LTL.
@@ V Problem Formulation
We first introduce a variant of the product team model defined in work [8], which aggregates the set of PBA for a given specification. Following this, we develop the hierarchical team model that integrates a set of product team models to address hierarchical sc-LTL specifications.
V-A Product Team Model
@@
Every leaf specification $\phi\in\Phi^{K}$ can be decomposed into a set of tasks such that
  • Independence: every task can be executed independently;
@@
The decomposition set ${\mathcal{D}}_{\mathcal{B}}\subseteq{\mathcal{Q}}_{\mathcal{B}}$ of the NBA ${\mathcal{B}}$ contains all states $q$ such that each state leads to a decomposition of tasks.
@@
Consider an sc-LTL specification $\phi=\Diamond a\wedge\Diamond b$, with its corresponding NBA depicted in Fig. 1(b). For any two-part satisfying word $\sigma=\sigma_{1}\sigma_{2}$, where $\sigma_{1}$ leads to a run ending at state 2, the reverse sequence $\sigma_{2}\sigma_{1}$ also satisfies the specification by leading to a run through state 1. This indicates that the tasks before and after state 2 can be completed in any order. Therefore, state 2 belongs to the decomposition set. Similarly, state 1 also belongs to the decomposition set. By default, all initial and accepting states in an NBA are considered decomposition states. $\square$
Figure 2: The product team model featuring two robots. PBAs ${\mathcal{P}}(1,\phi)$ and ${\mathcal{P}}(2,\phi)$ may be different reflecting the robots’ heterogeneity. Initial states are indicated with incoming arrows, and accepting states are depicted with double circles. The decomposition sets are highlighted in gray. In-spec switch transitions are represented by dashed arrows. A plan exclusive to robot 1 is traced by the green path, whereas the red path delineates a plan that involves the collaboration of both robots.
Definition V.3 (Product Team Model (modified from [8]))
Given a leaf specification $\phi\in\Phi^{K}$, the product team model ${\mathcal{P}}(\phi)$ consists of $N$ product models ${\mathcal{P}}(r,\phi)$ and is given by the tuple ${\mathcal{P}}(\phi):=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,s,q):r\in[N],(s,q)\in{\mathcal{Q}}_{\mathcal{P}}(r,\phi)\right\}$ is the set of state;
@@
  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation. One type of transition, referred to as in-spec transitions, occurs inside a PBA. That is, $(r,s,q)\to_{\mathcal{P}}(r,s^{\prime},q^{\prime})$ if $(s,q)\to_{{\mathcal{P}}(r,\phi)}(s^{\prime},q^{\prime})$; the other type of transition relation, denoted by $\zeta_{\text{in}}$, occurs between two PBAs and is defined below in Def. V.4;
  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)\}$ is the set of accepting/final states ;
@@
  • preserves the NBA progress, $q=q^{\prime}$;
  • represents a decomposition choice, $q\in{\mathcal{D}}_{\mathcal{B}}(\phi)$.
@@
Figure 3: Hierarchical team models for two leaf specifications and two robots. The product team models for each leaf specification are depicted within areas shaded in light orange and green. To enhance clarity, some transitions within a single PBA are omitted. Inter-spec type I and II switch transitions are illustrated using blue and red dashed arrows, respectively. It’s important to note that type I switch transitions are bidirectional, while type II switch transitions are unidirectional. For the sake of readability, inter-spec type I switch transitions are specifically shown for ${\mathcal{P}}(1,\phi_{1})$ and ${\mathcal{P}}(1,\phi_{2})$. Inter-spec type II switch transitions are only shown from ${\mathcal{P}}(2,\phi_{1})$ to ${\mathcal{P}}(1,\phi_{2})$. The path marked in yellow highlights a specific plan. This plan originates from the initial state of ${\mathcal{P}}(1,\phi_{1})$ and concludes at the accepting state of ${\mathcal{P}}(2,\phi_{2})$. This path involves assigning robot 2 to fulfill the specification $\phi_{1}$, which includes one in-spec switch transition ①. Subsequently, robots 1 and 2 are tasked with satisfying specification $\phi_{2}$, connected through a type II inter-spec switch transition ② and involving another in-spec switch transition ③.
V-B Hierarchical Team Models
To extend to hierarchical sc-LTL, we construct a product team model per leaf specification. The goal here is to connect them to form a connectable space. We refer to the scenario of decomposition of tasks for one specifications as in-spec independence. In contrast, leaf specifications can exhibit dependencies. For example, one leaf specification $\phi$ might need to be completed before another $\phi^{\prime}$. This scenario is referred to as inter-spec dependence. Another scenario is inter-spec independence, where two leaf specifications can be satisfied independently. We note that in-spec dependence does not exist, as precluded by Asm. V.1. To capture inter-spec dependency and independence, we introduce two additional types of switch transitions between different product team models.
@@ V-B Hierarchical Team Models
Given the hierarchical sc-LTL specifications $\Phi$, the hierarchical team models ${\mathcal{P}}$ consists of a set of product team models ${\mathcal{P}}(\phi)$ with $\phi\in\Phi^{K}$ and is denoted by the tuple ${\mathcal{P}}:=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,\phi,s,q):\phi\in\Phi^{K},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}(\phi)\right\}$ is the set of state;
  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:\phi\in\Phi^{K},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(\phi)\right\}$ is the set of initial states ;
  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation. One type of transition occurs inside a product team model, that is, $(r,\phi,s,q)\to_{\mathcal{P}}(r^{\prime},\phi,s^{\prime},q^{\prime})$ if $(r,s,q)\to_{{\mathcal{P}}(\phi)}(r^{\prime},s^{\prime},q^{\prime})$; another two types of transition relation, denoted by $\zeta_{\text{inter}}^{1}$ and $\zeta_{\text{inter}}^{2}$, occur between two product teams models and are defined below in Def. V.7 and V.8;
  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)\}$ is the set of accepting/final states.
@@
  • connects decomposition states, $q\in{\mathcal{D}}_{\mathcal{B}}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})$.
@@
  • connects an accepting state to a decomposition state, $q\in{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})$.
Inter-spec type I switch transitions enable switching to another PBA with a different leaf specification, while retaining the same robot and pausing the current task. Conversely, inter-spec type II switch transitions create a connection from the accepting state of one PBA to the decomposition state of another PBA associated with the first robot, indicating the completion of one specification and the resumption of another. These hierarchical team models can be graphically represented in a 3D space, with each product team model occupying a 2D plane; see Fig. 3. This structure has three axes: one representing the task progress for a specification (the $(s,q)$-axis), another indicating the task assignment among robots (the $r$-axis), and the last one tracking the overall progress of tasks across different specifications (the $\phi$-axis).
@@
The construction is not a product of all NBA and robot states. Instead, if each product team model is considered a distinct sub-space, the inter-spec switch transitions facilitate the connections between these sub-spaces. As a result, two consecutive product models are only loosely interconnected. The majority of states across different sub-spaces remain disconnected. Such a design significantly reduces the overall search space, enhancing the efficiency of the search.
Given a state-specification plan $\tau=\tau^{0}\tau^{1}\ldots\tau^{h}$ where $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$, the cost for robot $r$, such as energy consumption or completion time, is represented as $c_{r}=\sum_{i=0}^{h-1}c_{r}(s^{i}_{r},s^{i+1}_{r})$, where $c_{r}(s^{i}_{r},s^{i+1}_{r})$ denotes the cost incurred transitioning between states for robot $r$. The goal is to minimize the additive cost, expressed as
@@
Given the hierarchical team models ${\mathcal{P}}$, derived from transition systems of robots and the NBAs of hierarchical sc-LTL specifications $\Phi$, find an optimal state-specification plan $\tau^{*}$ that minimizes $J(\tau^{*})$ and satisfies $\Phi$.
VI Planning under Hierarchical sc-LTL

### v4

@@ 16             $\sigma^{\prime}_{j}=\varnothing$ ;
The syntax and semantics for H-LTLf here are inspired by the domain of robots and may not be applicable to other domains. Our experience indicates that a top-down reasoning approach is effective in constructing H-LTLf specifications from task descriptions. This entails recognizing the inclusion relationships among tasks and organizing them into various levels of abstraction, ranging from general to specific. At each level, tasks that are closely related should be combined into single formulas, while those with weaker connections should be grouped into separate formulas.
@@
The hierarchical format can be applied to syntactically co-safe LTL (sc-LTL), a syntactic subset of LTL, to formulate hierarchical sc-LTL. Sc-LTL formulas can be satisfied by finite sequences followed by any infinite repetitions.
@@ IV-C Expressiveness
Temporal logics are interpreted over paths in finite Kripke structures. The Kripke structure is identical to the transition system as presented in Def. III.2 in this work. A Kripke structure, denoted by ${\mathcal{K}}$, is deemed finite when it has a finite number of states. ${\mathcal{K}}$ is considered a model for a temporal logic formula $\phi$ if every initial finite path within ${\mathcal{K}}$ satisfies $\phi$.
@@
H-LTLf is more expressive than the standard LTLf.
@@
We first show that, for any standard LTLf, there exists an equivalent hierarchical counterpart. Therefore, the standard LTLf is subsumed by the H-LTLf. Given a standard LTLf $\phi$, an equivalent two-level H-LTLf can be constructed in a straightforward manner, where the top level is $\Diamond\phi$ and the bottom level is $\phi$ itself, that is,
@@
Next, we introduce a counter-example to demonstrate the absence of the standard equivalent for certain H-LTLf specifications. Considering a Kripke structure with two atomic propositions $\{a,b\}$, the following H-LTLf specifications include two mutually exclusive leaf specifications:
@@
$L_{2}:\quad$ $\phi_{2}^{1}=\neg b\,\mathcal{U}\,(a\wedge\neg b),\quad\phi_{2}^{2}=\neg a\,\mathcal{U}\,(b\wedge\neg a).$
While leaf specifications are mutually exclusive, the introduction of the state-specification sequence enables the association of each state with one specification, facilitating the separate treatment of leaf specifications. For example, when these two specifications are associated with different Kripke structures, they can be realistically fulfilled. However, when solely using atomic propositions $a$ and $b$, combining two leaf specifications into an LTLf specification is unfeasible, as it inevitably results in a contradiction, reducing the specification to False. ∎
@@ V Problem Formulation
We first present an assumption on H-LTLf, and then introduce a variant of the product team model defined in work [9], which aggregates the set of PA for a given specification. Following this, we develop the hierarchical team model that integrates a set of product team models.
V-A Assumptions on H-LTLf
When considering the composite propositions in the context of robot tasks, they correspond to sub-tasks. For example, one such task could be retrieving a beverage from the refrigerator, such as a bottle of apple juice or orange juice, where completing just one sub-task suffices. Based on this observation, we assume that, for a task to be successfully completed, it is sufficient that each of its sub-tasks occurs at most once. This assumption applies only to non-leaf specifications.
Assumption V.1 (Bounded One-Time Satisfaction)
For any non-leaf specification $\phi_{k}^{i}$ at level $L_{k}$, there exists an accepting input word $\sigma=\sigma_{0},\sigma_{1},\dots,\sigma_{n}$ where $\sigma_{j}\subseteq\texttt{Prop}(\phi_{k}^{i})\subseteq\Phi_{k+1}$ for $j\in[n]$, such that $\sum_{j=0}^{n}\mathbb{I}(\phi_{k+1}^{l}\in\sigma_{j})\leq 1,$ for any $l\in[|\Phi_{k+1}|]_{+}$, where $\mathbb{I}$ is the indicator function.
V-B Product Team Model
@@
Every leaf specification $\phi\in\Phi_{\text{leaf}}$ can be decomposed into a set of tasks such that
  • Independence: tasks can be executed independently;
@@
The decomposition set ${\mathcal{D}}_{\mathcal{A}}\subseteq{\mathcal{Q}}_{\mathcal{A}}$ of the NFA ${\mathcal{A}}$ contains all states $q$ such that each state leads to a decomposition of tasks.
@@
Consider an LTLf specification $\phi=\Diamond a\wedge\Diamond b$, with its corresponding NFA depicted in Fig. 1(b). For any two-part satisfying word $\sigma=\sigma_{1}\sigma_{2}$, where $\sigma_{1}$ leads to a run ending at state 2, the reverse sequence $\sigma_{2}\sigma_{1}$ also satisfies the specification by leading to a run through state 1. This indicates that the tasks before and after state 2 can be completed in any order. Therefore, state 2 belongs to the decomposition set. Similarly, state 1 also belongs to the decomposition set. By default, all initial and accepting states in an NFA are considered decomposition states. $\square$
Definition V.4 (Product Team Model [9])
Given a leaf specification $\phi\in\Phi_{\text{leaf}}$, the product team model ${\mathcal{P}}(\phi)$ consists of $N$ product models ${\mathcal{P}}(r,\phi)$ and is given by the tuple ${\mathcal{P}}(\phi):=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,s,q):r\in[N]_{+},(s,q)\in{\mathcal{Q}}_{\mathcal{P}}(r,\phi)\right\}$ is the set of state;
@@
  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation. One type of transition, referred to as in-spec transitions, occurs inside a PA. That is, $(r,s,q)\to_{\mathcal{P}}(r,s^{\prime},q^{\prime})$ if $(s,q)\to_{{\mathcal{P}}(r,\phi)}(s^{\prime},q^{\prime})$; the other type of transition relation, denoted by $\zeta_{\text{in}}$, occurs between two PAs and is defined below in Def. V.5;
  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)\}$ is the set of accepting/final states ;
@@
  • preserves the NFA progress, $q=q^{\prime}$;
  • represents a decomposition choice, $q\in{\mathcal{D}}_{\mathcal{A}}(\phi)$.
Figure 3: The product team model featuring two robots. PAs ${\mathcal{P}}(1,\phi)$ and ${\mathcal{P}}(2,\phi)$ may be different reflecting the robots’ heterogeneity. Initial states are indicated with incoming arrows, and accepting states are depicted with double circles. The decomposition sets are highlighted in gray. In-spec switch transitions are represented by dashed arrows. A plan exclusive to robot 1 is traced by the green path, whereas the red path delineates a plan that involves the collaboration of both robots.
@@
Figure 4: Hierarchical team models for two leaf specifications and two robots. The product team models for each leaf specification are depicted within areas shaded in light orange and green. To enhance clarity, some transitions within a single PA are omitted. Inter-spec type I and II switch transitions are illustrated using blue and red dashed arrows, respectively. It’s important to note that type I switch transitions are bidirectional, while type II switch transitions are unidirectional. For the sake of readability, inter-spec type I switch transitions are specifically shown for ${\mathcal{P}}(1,\phi_{1})$ and ${\mathcal{P}}(1,\phi_{2})$. Inter-spec type II switch transitions are only shown from ${\mathcal{P}}(2,\phi_{1})$ to ${\mathcal{P}}(1,\phi_{2})$. The path marked in yellow highlights a specific plan. This plan originates from the initial state of ${\mathcal{P}}(1,\phi_{1})$ and concludes at the accepting state of ${\mathcal{P}}(2,\phi_{2})$. This path involves assigning robot 2 to fulfill the specification $\phi_{1}$, which includes one in-spec switch transition ①. Subsequently, robots 1 and 2 are tasked with satisfying specification $\phi_{2}$, connected through a type II inter-spec switch transition ② and involving another in-spec switch transition ③.
V-C Hierarchical Team Models
To extend to H-LTLf, we construct a product team model per leaf specification. The goal here is to connect them to form a connectable space. We refer to the scenario of decomposition of tasks for one specifications as in-spec independence. In contrast, leaf specifications can exhibit dependencies. For example, one leaf specification $\phi$ might need to be completed before another $\phi^{\prime}$. This scenario is referred to as inter-spec dependence. Another scenario is inter-spec independence, where two leaf specifications can be satisfied independently. We note that in-spec dependence does not exist, as precluded by Asm. V.2. To capture inter-spec dependency and independence, we introduce two additional types of switch transitions between different product team models.
@@ V-C Hierarchical Team Models
Given the H-LTLf specifications $\Phi$, the hierarchical team models ${\mathcal{P}}$ consists of a set of product team models ${\mathcal{P}}(\phi)$ with $\phi\in\Phi_{\text{leaf}}$ and is denoted by the tuple ${\mathcal{P}}:=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,\phi,s,q):\phi\in\Phi_{\text{leaf}},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}(\phi)\right\}$ is the set of state;
  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:\phi\in\Phi_{\text{leaf}},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(\phi)\right\}$ is the set of initial states;
  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation. One type of transition occurs inside a product team model, that is, $(r,\phi,s,q)\to_{\mathcal{P}}(r^{\prime},\phi,s^{\prime},q^{\prime})$ if $(r,s,q)\to_{{\mathcal{P}}(\phi)}(r^{\prime},s^{\prime},q^{\prime})$; another two types of transition relation, denoted by $\zeta_{\text{inter}}^{1}$ and $\zeta_{\text{inter}}^{2}$, occur between two product teams models and are defined below in Def. V.8 and V.9;
  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)\}$ is the set of accepting/final states.
@@
  • connects decomposition states, $q\in{\mathcal{D}}_{\mathcal{A}}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})$.
@@
  • connects an accepting state to a decomposition state, $q\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})$.
Inter-spec type I switch transitions enable switching to another PA with a different leaf specification, while retaining the same robot and pausing the current task. Conversely, inter-spec type II switch transitions create a connection from the accepting state of one PA to the decomposition state of another PA associated with the first robot, indicating the completion of one specification and the resumption of another. It points toward the first robot, the starting robot of a product team model, because the in-spec switch transition is unidirectional, always pointing to the next robot. These hierarchical team models can be graphically represented in a 3D space, with each product team model occupying a 2D plane; see Fig. 4. This structure has three axes: one representing the task progress for a specification (the $(s,q)$-axis), another indicating the task assignment among robots (the $r$-axis), and the last one tracking the progress of different specifications (the $\phi$-axis).
@@
The construction is not a product of all NFA and robot states. Instead, if each product team model is considered a distinct sub-space, the inter-spec switch transitions facilitate the connections between these sub-spaces. As a result, two consecutive product models are only loosely interconnected. The majority of states across different sub-spaces remain disconnected. Such a design significantly reduces the overall search space, enhancing the efficiency of the search.
Given a state-specification sequence $\tau=\tau_{0}\tau_{1}\ldots\tau_{h}$ where $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),(s_{2}^{i},\psi_{2}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$, the cost for robot $r$, such as energy consumption or completion time, is represented as $c_{r}=\sum_{i=0}^{h-1}c_{r}(s^{i}_{r},s^{i+1}_{r})$, where $c_{r}(s^{i}_{r},s^{i+1}_{r})$ denotes the cost incurred transitioning between states for robot $r$. The goal is to minimize the additive cost, expressed as
@@
Given the hierarchical team models ${\mathcal{P}}$, derived from transition systems and the H-LTLf specifications $\Phi$ that satisfy Asm. V.1, find an optimal state-specification sequence $\tau^{*}$ that satisfies $\Phi$ and minimizes $J(\tau^{*})$.
VI Planning under H-LTLf